A 25.0 L helium balloon was filled in a desert at temperature of 35 oC. When the balloon is released to 20,000 ft, the air temperature becomes -48 oC. What is the new volume if the atmosphere pressure remains unchanged? Answer in 3 sig-dig number without unit L


1
Expert's answer
2021-03-22T06:32:46-0400

"\frac{P_1V_1}{T_1}" = "\frac{P_2V_2}{T_2}"

Pressure is unchanged

"\frac{V_1}{T_1} = \frac{V_2}{T_2}"


"\frac{25.0L}{35⁰C} = \frac{V/2}{-48⁰C}"


V2 = "\frac{25×(-48)}{35}" = -34.285 L


V2 = 34.285 L


3 sig digit


V2 = 342.85 × 10-1 L